What annual rate of return is earned on a $4,000 investment made in year 5 when it grows to $8,200 by the end of year 8? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Annual Rate of Return = [(Ending Value / Begining Value) ^ (1 / number of years)] - 1 = [(8,200 / 4000) ^ (1 / 3)] - 1 = 1.27 - 1 = 0.27 = 27%